OpenStudy (anonymous):

Calculate the derivative of the function. f(x) = (x+8)^-2

OpenStudy (anonymous):

i tried (x+8)^-3 already

OpenStudy (anonymous):

f(x) = (x+8)^-2 f'(x)=(-2)(x+8)^-3

OpenStudy (anonymous):

This is just the regular differentiation rule: The derivative of x^(-2) is -2*x^(-3). In the same way, and as CoCo said, the derivative of (x+c)^(-2) is -2*(x+c)^(-3) for any constant c. This is easy to see by the u substitution u = x+c.

OpenStudy (anonymous):

got it